State the formula for sin(a + b) and use it to show sin(a + 3π/2) = − cos(a)
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

State the formula for sin(a + b) and use it to show sin(a + 3π/2) = − cos(a)
---------
sin(a + b) = sin(a)cos(b) + cos(a)sin(b)
----------
sin(a+(3pi/2)) sin(a)cos(3pi/2) + cos(a)(sin(3pi/2))
-------
= (sin(a))*0 + cos(a)(-1)
-------
= 0 - cos(a)
-------
= -cos(a)
================
Cheers,
Stan H.
================

Answer by lwsshak3(11628)

State the formula for sin(a + b) and use it to show sin(a + 3π/2) = − cos(a)
sin(a+b)=sina*cosb+cosa*sinb
sin(a+3π/2)=sina*0+cosa*-1=-cos(a)